OpenStudy (anonymous):

convert the polar equation to rectangular form. r=6/(2-3sin theta)

OpenStudy (anonymous):

Can anyone help?

OpenStudy (anonymous):

ok in order to convert use the following \[\Large x^2+y^2=r^2\] and \[\Large y=rsin(\theta)\]

OpenStudy (anonymous):

can you do this?

OpenStudy (anonymous):

Sort of.

OpenStudy (anonymous):

ok replace sin(theta)=y/r so \[\Large (\sqrt{x^2+y^2})=\frac{6}{2-3y/r}\]

OpenStudy (anonymous):

use again \[\Large r=\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

is it ok now ?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

yw:)